OpenStudy (anonymous):

Find the constant of variation for the equation z=1/3x

OpenStudy (anonymous):

A ) z B) x C) 1/3 D) no constant

OpenStudy (anonymous):

@phi @inkyvoyd @Yacoub1993 @greenlegodude57 @francoisderozay @djcool31 @e.mccormick @arabpride @whpalmer4 @quixoticideals @Vallarylynn16 @greenlegodude57 @mlb004 @djcool31

OpenStudy (whpalmer4):

Are we finding the history of this equation? :-) Is the equation supposed to be \[z = \frac{1}{3}z\] or \[z = \frac{1}{3z}\]?

OpenStudy (whpalmer4):

Assuming it is the first one: Direct variation between \(z\) and \(x\) is given by the form\[z=kx\] where \(k\) is the constant of variation.

OpenStudy (whpalmer4):

If it's the second one, that's an indirect variation between \(z\) and \(x\) and is given by the form\[z = \frac{k}{x}\]where, again, \(k\) is the constant of variation.

OpenStudy (anonymous):

@whpalmer4 idk those are the choices they gave me

OpenStudy (whpalmer4):

What don't you know?

OpenStudy (anonymous):

so would it be Z

OpenStudy (whpalmer4):

Is \(z\) a constant? Why would you choose \(z\)?

OpenStudy (whpalmer4):

Did you read my posts at all? Did any of them refer to \(z\) as the constant of variation?

OpenStudy (anonymous):

OHH SO ITS NO CONSTANT

OpenStudy (whpalmer4):

Come on. First, which of the two equations I showed most closely resembles the one in your problem?

OpenStudy (anonymous):

1/3 ?

OpenStudy (whpalmer4):

Will you please answer my question? Which of the two formulas here most closely resembles what is on your screen? \[z = \frac{1}{3}x\]or\[z=\frac{1}{3x}\]

OpenStudy (anonymous):

THE FIRST

OpenStudy (whpalmer4):

Excellent. That means it is direct variation. Direct variation between \(x\) and \(z\) is written in the form \[z = kx\]where \(k\) is a constant called the constant of variation. Compare your equation with mine: \[z = \frac{1}{3}x\]\[z=kx\] What is the constant of variation?